OpenStudy (anonymous):

A car moves around a curve whose radius is 320m. The car's speed is 28m/s. What is the coefficient of static friction between the road and the tires? Oh and the centripetal force is 2.45m/s/s

OpenStudy (kainui):

So the equation for centripetal force has the same units as acceleration, \[a_c=\frac{ v^2 }{ r }\] Now the formulas for force are: F=ma and F=un where u is the coefficient of static friction and n is the normal force. Now we can brew these altogether and see that the forces are in equilibrium: \[m \frac{ v^2 }{ r }=umg\] And hey, what do you know, we can solve for u without actually knowing the mass of the car!